OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (2, 5) and (3, 0). y − = (x − 2)

OpenStudy (igreen):

First we find the slope. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{0-5}{3-2}\) Subtract & Divide

OpenStudy (anonymous):

\[\frac{ -5 }{ 1 } = -5\]

OpenStudy (igreen):

Yes, so we can plug -5 into the equation. \(y - y_1 = m(x- x_1)\) Where 'm' is slope: \(y - y_1 = -5(x - x_1)\) Now we can plug any of the two points into this equation. \(x_1\) is the x-value of the point, and \(y_1\) is the y-value of the point.

OpenStudy (igreen):

So can you plug in (2, 5)?

OpenStudy (anonymous):

\[y-5=-5(x-2)\]

OpenStudy (igreen):

Yes!

OpenStudy (igreen):

That's your answer.

OpenStudy (anonymous):

thank you for the help..again lol

OpenStudy (igreen):

No problem :)